# Optimize the Unseen - Fast NeRF Cleanup with Free Space Prior

### v1

@@ Abstract
Neural Radiance Fields (NeRF) have advanced photorealistic novel view synthesis, but their reliance on photometric reconstruction introduces artifacts, commonly known as ”floaters”. These artifacts degrade novel view quality, especially in areas unseen by the training cameras. We present a fast, post-hoc NeRF cleanup method that eliminates such artifacts by enforcing our Free Space Prior, effectively minimizing floaters without disrupting the NeRF’s representation of observed regions. Unlike existing approaches that rely on either Maximum Likelihood (ML) estimation to fit the data or a complex, local data-driven prior, our method adopts a Maximum-a-Posteriori (MAP) approach, selecting the optimal model parameters under a simple global prior assumption that unseen regions should remain empty. This enables our method to clean artifacts in both seen and unseen areas, enhancing novel view quality even in challenging scene regions. Our method is comparable with existing NeRF cleanup models while being $\mathbf{2.5\times}$ faster in inference time, requires no additional memory beyond the original NeRF, and achieves cleanup training in less than 30 seconds. Our code will be made publically available.
@@ 1 Introduction
Figure 1: Density Cleanup: Areas unseen by training rays lack optimization, leading to artifacts and ”floaters”. To demonstrate this, we sampled densities uniformly within the bounding box [-1, 1], generating a point cloud of regions with densities above a threshold. Training cameras marked in red and evaluation cameras marked in green. The Top scene highlights significant floater artifacts, especially behind training cameras where rays are absent, which can degrade novel view quality. Our approach, as shown on the bottom, effectively removes these artifacts by optimizing densities even in unseen regions, resulting in cleaner novel views.
@@
Neural Radiance Fields (NeRF) have emerged as a leading technique in photorealistic scene reconstruction and novel view synthesis, effectively capturing complex scenes from a limited set of images. However, NeRF’s reliance on photometric optimization introduces a significant limitation: it performs poorly in regions of the scene that lack direct observations from the training images. This results in visual artifacts, commonly referred to as ”floaters,” which are density accumulations in areas unseen by the training cameras. These floaters degrade rendering quality, especially in novel views where the artifacts may obscure scene surfaces or introduce false details, as illustrated in Figure 1.
@@
While state-of-the-art approaches for NeRF cleanup can mitigate artifacts, they often come with a high computational cost. Nerfbusters [26] uses 3D diffusion models to generate a learned, data-driven 3D prior, necessitating NeRF fine-tuning and resulting in substantial computational demands and extended cleanup training time. BayesRays [8], an alternative approach, reduces training time by performing post-hoc uncertainty-based cleanup; however, it introduces considerable overhead at inference due to continuous uncertainty estimation. These approaches highlight a key trade-off in existing methods between efficient artifact removal and computational expense, underscoring the need for a more effective solution.
Current methods follow either a Maximum Likelihood (ML) approach, optimizing model parameters to fit the observed data, or rely on a complex, data-driven local prior to match a general data distribution. We, in contrast, take a Maximum-a-Posteriori (MAP) approach with a simple global prior. Where, among all possible models that can explain the data, we seek the one that also satisfies the prior assumption.
We introduce a novel, post-hoc method that cleans up artifacts in pre-trained NeRFs without altering the core NeRF architecture or adding any extra computation during inference. Similarly to Gaussian Splatting [11], that ideally models only visible regions by assuming unobserved regions to be empty, our method applies this assumption to NeRF. In practice, however, Gaussian Splatting can still suffer from artifacts, and NeRF is even more prone to unwanted density accumulations (”floaters”) in unseen areas. By explicitly optimizing these regions in NeRF, our method reduces floaters and improves scene quality, outperforming even Gaussian Splatting in cleanup.
@@
Our method samples across the entire 3D space, applying a Free Space Prior that enforces zero density in regions unobserved by training rays. Our method is designed to be lightweight and efficient: we only fine-tune the NeRF itself, without training additional networks or increasing inference-time memory and computation. This ensures that floaters are minimized without disrupting the scene’s intended structure or introducing additional memory requirements. The method is both efficient and highly compatible with existing NeRF cleanup models, offering a significant improvement in speed and memory efficiency compared to state-of-the-art artifact removal techniques. The contributions of our method are summarized as follows:
@@
  • Efficiency in computation and memory: Our approach requires no additional memory beyond the original NeRF, is 2.5x faster in inference time, and 1.5x faster in cleanup train time to current SOTA methods.
  • State-of-the-art artifact removal: Our method achieves high-quality NeRF cleanup by effectively eliminating floaters, leading to more accurate novel view synthesis.
@@
  • Optimization in unseen regions: Unlike previous methods that focus solely on regions observed by training cameras, our method also optimizes density in unseen regions, reducing artifacts throughout the scene.
2 Related Works
Neural Radiance Fields (NeRFs) [17] often exhibit visual artifacts, such as ”floaters” and density inconsistencies, due to training data imperfections, including incomplete scene coverage and data noise. These issues introduce uncertainties in unobserved regions, which can be divided into aleatoric and epistemic uncertainty.
Aleatoric uncertainty, arising from data noise like lighting changes or transient objects, has been addressed in dynamic scenes by works such as NeRF-W [13] and others [21, 20, 9]. In contrast, our focus is on static scenes with epistemic uncertainty, stemming from incomplete scene information. Approaches like CF-NeRF [23] use variational inference to capture this uncertainty in geometry, while BayesRays [8] applies a Bayesian framework to estimate epistemic uncertainty post-training. Our work specifically addresses cleanup of artifacts in static scenes caused by epistemic uncertainty.
Figure 2: Architecture Overview: Our method fine-tunes a pre-trained NeRF to clean artifacts. During each optimization step, we simultaneously optimize two objectives: (1) density is optimized using our Free Space Prior across randomly sampled 3D points, enforcing zero density in all regions, and (2) the NeRF reconstruction loss is optimized along the original training images. For clarity, the points are grouped by region: blue points represent areas unseen by the training cameras, while green and red points are seen by the training cameras. Green points lie between the camera and the object, while red points are located on the object surface. Note that blue points may also appear along camera rays due to occlusions.
@@ 2 Related Works
Various methods have been proposed to address the visual artifacts common in NeRF reconstructions. For example, Nerfbusters [26] introduced a post-processing technique using diffusion models that learns a local 3D data-driven prior from synthetic data. This data-driven prior, incorporated during NeRF optimization, encourages plausible geometry but requires extensive training and NeRF fine-tuning, adding significant computational overhead. Another approach, BayesRays [8], leverages a post-hoc solution by thresholding an uncertainty field to mask high-uncertainty regions, effectively cleaning artifacts at inference time but at the cost of increased computation during rendering.
These methods, though effective, often come with trade-offs in memory consumption, training time, or inference speed. In contrast, our approach removes artifacts efficiently, preserving the original NeRF’s structure and performance without requiring additional models or extensive post-processing, based only on our Free Space Prior. This ensures both speed and practicality in artifact cleanup.
@@
Data-driven optimization is standard in deep learning, but numerous studies demonstrate that incorporating a well-chosen prior can significantly boost performance. For instance, priors have proven effective in inverse problems like image denoising [4, 10, 28, 22, 5, 6]. In particular, [5, 6] introduce a global image prior enforcing sparsity, showcasing how a Bayesian framework can yield a simple yet powerful denoising algorithm. In the context of 3D reconstruction, priors and regularization play similarly critical roles. Nerfbusters [26] uses a local prior to improve scene fidelity and and RegNeRF [19] employs a 2D Total Variation (TV) regularizer to rendered images, both maintaining the standard NeRF architecture as we do. Other methods modify the architecture. For example, Plenoxels [30] employs a 3D Total Variation (TV) regularizer to reduce abrupt density changes between neighboring voxels and 2D TV regularization is also applied in factorized plenoptic fields [3, 7]. Plenoctrees [31] incorporates a sparsity loss within a complex pipeline for faster rendering but at high memory and training costs. Alternatively, mip-NeRF 360 [2] applies a distortion loss to penalize regions with high photometric error, offering regularization that reduces artifacts and enhances model consistency.
@@ 3 Background
We propose a method for cleaning artifacts in Neural Radiance Fields (NeRF) by addressing the limitations of conventional photometric optimization. Our approach leverages the strengths of photometric optimization to preserve the scene’s integrity while effectively eliminating artifacts. Understanding NeRF is essential to grasp the significance of our work and its contributions to the field. In this section, we review the fundamental concepts of NeRF and discuss the challenges associated with artifact generation in unobserved regions, which our approach aims to overcome.
@@ 3.1 Neural Radiance Fields
Neural Radiance Fields (NeRF) [17] represent an approach for synthesizing novel views of complex 3D scenes. In NeRF, each point in 3D space is represented by a view-dependent radiance and a view-independent density value, mathematically expressed as follows:
@@
Here, $\boldsymbol{\phi}$ denotes the learnable parameters of the neural network. The color of each pixel in the rendered image is computed through volume rendering techniques, where the color along a ray $\mathbf{r}=\mathbf{o}_{\mathbf{r}}+t\cdot\mathbf{d}_{\mathbf{r}}$ is determined by sampling points ${t_{i}}$ along the ray:
@@
An important observation from Equation 2 is that NeRF’s optimization occurs on a per-pixel basis, with the color of each pixel computed along a single ray. Since density decreases exponentially along the ray after it intersects the first surface, there is effectively no optimization beyond that surface, and certainly none where the ray does not pass through. Consequently, unseen regions remain unoptimized, often resulting in unwanted artifacts in these areas.
@@ 4 Method
Our goal is to eliminate floater artifacts in NeRF reconstructions while preserving the structural integrity of the scene. Our approach, as shown in our architecture Figure 2, is designed as a post-hoc refinement, takes a pre-trained NeRF model along with its training cameras and optimizes it to remove artifacts in unseen regions while retaining the scene’s intended features.
To date, NeRF models, and NeRF cleanup methods in particular, searched for a density field $\sigma$ that best explains the given data. That is, They look for the maximum likelihood solution to the problem. Some methods extend this by learning a general data distribution over multiple 3D scenes, creating a local prior for artifact removal. In contrast, we apply a simple, global prior on the density $\sigma$ We take the prior $P(\sigma)$ to be the zero density prior.
Optimizing the Unseen
NeRF excels at minimizing photometric loss over the training images, providing accurate reconstructions in observed areas. However, it does not constrain unseen regions, often resulting in noisy artifacts, or ”floaters,” in those areas. We base our method on the prior assumption that “There is nothing in the unseen regions.” While this assumption does not necessarily hold in real-world scenarios, for NeRF reconstructions, it is more practical than allowing unseen regions to contain unregulated noise. To enforce this assumption, we apply a sigmoid-softened density prior term that gently drives density $\sigma$ toward zero:
$sigmoid(\sigma)=\frac{1}{1+e^{-\sigma}}$ (4)
@@
Ideally, we would like to directly set densities to zero only in the unseen regions. However, determining which regions are strictly unobserved by training rays is challenging due to scene occlusions and structural complexity. Additionally, even if feasible, this direct approach would be computationally expensive. rather than explicitly setting densities, we apply the Free Space Prior uniformly across the entire 3D space, affecting both seen and unseen regions. This approach simplifies computation while guiding NeRF optimization to achieve cleaner density distributions.
To implement this, we randomly sample $\mathcal{N}$ points across the 3D space and query the NeRF model $\boldsymbol{\phi}$ for densities at these points, resulting in a set $\Sigma=\{\sigma(\mathbf{x}_{i})\mid\mathbf{x}_{i}\in\mathbb{R}^{3},\,i=1,\dots,\mathcal{N}\}$ of density values. We then construct a softened density term for the Free Space Prior, as shown in Equation 4, which we integrate into a Free Space Prior Loss term:
$\mathcal{L}_{\text{FSP}}=\sum_{i=1}^{\mathcal{N}}\left(\frac{1}{1+e^{-\sigma(\mathbf{x}_{i})}}\right)^{2}$ (5)
This Free Space Prior Loss optimizes the NeRF parameters $\boldsymbol{\phi}$ by enforcing densities to be zero, reflecting our assumption that “There is nothing in the unseen regions.” However, this loss also inadvertently reduces densities in seen regions, where densities should ideally reflect the actual scene structure.
Figure 3: Qualitative Cleanup Results: Visualization of different cleanup methods applied to the Plant and Roses scenes from the Nerfbusters dataset. Each image shows a novel view from the evaluation set rendered post-cleanup. The comparison highlights each method’s effectiveness in balancing artifact removal with scene coverage. These visual results correspond to Figure 4 and illustrate the concept of coverage: methods such as Free Space Prior and BayesRays with a 0.9 threshold yield high coverage with minimal artifacts, while Nerfbusters and BayesRays with a 0.3 threshold result in lower coverage, with varying degrees of artifact cleanup. For comparison, we also provide in the last column the results of Gaussian Splatting.
Clean While Preserve
Applying the Free Space Prior Loss indiscriminately across 3D space can lead to an empty scene, as the randomly enforced zero-density constraint propagates, causing the MLP to approximate densities as zero throughout the space. To counter this, we introduce a combination of two competing losses: the Free Space Prior Loss Equation 5 to reduce densities in unseen regions and the NeRF photometric loss Equation 3 applied along training rays to preserve the scene structure and appearance.
@@
  • Unseen Regions: Only the Free Space Prior Loss affects these regions, pushing densities toward zero and removing floaters.
@@
The overall loss function for our method is a weighted combination of these two terms, where $\lambda$ acts as a temperature parameter, controlling the aggressiveness of the cleanup process.
$\mathcal{L}=\mathcal{L}_{\text{rec}}+\lambda\mathcal{L}_{\text{prior}}$ (6)
To summarize, our method provides an efficient solution for NeRF artifact cleanup by applying a Free Space Prior to unseen regions, effectively eliminating floaters without adding extra networks or computations during inference. By fine-tuning the pre-trained NeRF, we achieve significant cleanup with minimal computational overhead. The combination of the Free Space Prior loss and photometric loss allows for effective removal of artifacts while preserving the scene structure, making our method faster and more memory-efficient compared to existing approaches. Our experiments demonstrate that this approach maintains high-quality scene reconstruction with reduced training and inference times.

### v3

@@ Abstract
Neural Radiance Fields (NeRF) have advanced photorealistic novel view synthesis, but their reliance on photometric reconstruction introduces artifacts, commonly known as "floaters". These artifacts degrade novel view quality, particularly in unseen regions where NeRF optimization is unconstrained. We propose a fast, post-hoc NeRF cleanup method that eliminates such artifacts by enforcing a Free Space Prior, ensuring that unseen regions remain empty while preserving the structure of observed areas. Unlike existing approaches that rely on Maximum Likelihood (ML) estimation or complex, data-driven priors, our method adopts a Maximum-a-Posteriori (MAP) approach with a simple yet effective global prior. This enables our method to clean artifacts in both seen and unseen areas, significantly improving novel view quality even in challenging scene regions. Our approach generalizes across diverse NeRF architectures and datasets while requiring no additional memory beyond the original NeRF. Compared to state-of-the-art cleanup methods, our method is $\mathbf{2.5\times}$ faster in inference and completes cleanup training in under 30 seconds.
@@ 1 Introduction
Figure 1: Density Cleanup Across NeRF Variants and Datasets: The top row shows novel views taken near the training cameras, while the middle row presents novel views from farther away. The bottom row demonstrates the same distant views after applying our cleanup method, effectively removing artifacts while preserving scene integrity. Our approach generalizes well across a wide range of NeRF methods and datasets - see supplementary material for video demonstrations.
@@
Neural Radiance Fields (NeRF) have emerged as a leading technique in photorealistic scene reconstruction and novel view synthesis, effectively capturing complex scenes from a limited set of images. However, NeRF’s reliance on photometric optimization introduces a significant limitation: it performs poorly in regions of the scene that lack direct observations from the training images. This results in visual artifacts, commonly referred to as "floaters", which are density accumulations in areas unseen by the training cameras. These floaters degrade rendering quality, especially in novel views where the artifacts may obscure scene surfaces or introduce false details, as illustrated in Fig.˜1.
@@
While state-of-the-art approaches for NeRF cleanup can mitigate artifacts, they often come with a high computational cost. Nerfbusters [31] uses 3D diffusion models to generate a learned, data-driven 3D prior. This requires NeRF fine-tuning and results in substantial computational demands and extended cleanup training time. BayesRays [10], an alternative approach, reduces cleanup time by performing post-hoc uncertainty-based cleanup; however, it does not modify the NeRF scene itself. Instead, it applies uncertainty-based filtering during inference to clean the rendered frames, which introduces considerable overhead as each frame requires additional uncertainty estimation and processing. These approaches highlight a key trade-off in existing methods between efficient artifact removal and computational expense, underscoring the need for a more effective solution.
Unlike previous methods, which often require architecture-specific modifications or extensive fine-tuning, our approach is robust and generalizes effectively across diverse NeRF architectures and datasets, as demonstrated in Fig.˜1. This flexibility makes our method applicable to a wide range of scenarios, outperforming other cleanup methods that are constrained by their reliance on specific priors or additional computational overhead.
@@
A key contribution of our work is that we sample across the entire 3D space, not just along rays. Samples along the rays ensure NeRF fits the given scene, while samples in 3D space introduce a Free Space Prior. This contribution makes our approach lightweight and efficient: we only fine-tune the NeRF itself, without training additional networks or increasing inference-time memory and computation. As a result, floaters are minimized without disrupting the scene’s intended structure or introducing additional memory requirements. The method is both efficient and highly compatible with existing NeRF cleanup models, offering a significant improvement in speed and memory efficiency compared to state-of-the-art artifact removal techniques.
The contributions of our method are summarized as follows:
@@
  • Efficient and Lightweight: Our approach requires no additional memory beyond the original NeRF, is 2.5× faster in inference time, and 1.5× faster in cleanup training compared to current state-of-the-art methods.
@@
  • Optimization in Unseen Regions: Unlike previous methods that refine only regions observed by training cameras, our method actively optimizes density across unseen regions, enforcing a clean 3D representation and reducing artifacts throughout the scene.
  • Robust and Generalizable: Our approach generalizes seamlessly across different NeRF architectures and datasets, requiring no architectural modifications or learned priors, making it widely applicable to various real-world scenarios.
  • State-of-the-Art Artifact Removal: Our method effectively eliminates floaters and other density artifacts, enhancing novel view synthesis quality while preserving scene details.
2 Related Work
Figure 2: Overview of Our Cleanup Method: Our approach fine-tunes a pre-trained NeRF by optimizing density across both seen and unseen regions. We sample points in two ways: (1) along the original training rays (green points), maintaining consistency with the observed scene structure, and (2) randomly across the entire 3D space (blue points), enforcing our Free Space Prior to remove artifacts in unseen areas. Importantly, we also apply the Free Space Prior behind scene surfaces, ensuring that empty regions remain free of unwanted density accumulations. This ensures effective cleanup of floaters while preserving scene integrity.
@@ 2 Related Work
Neural Radiance Fields (NeRFs) [22] often exhibit visual artifacts, such as "floaters" and density inconsistencies, due to training data imperfections, including incomplete scene coverage and data noise. These issues introduce uncertainties in unobserved regions, which can be divided into aleatoric and epistemic uncertainty. Aleatoric uncertainty, arising from data noise like lighting changes or transient objects, has been addressed in dynamic scenes by works such as NeRF-W [18] and others [26, 25, 11]. In contrast, our focus is on static scenes with epistemic uncertainty, stemming from incomplete scene information. Approaches like CF-NeRF [28] use variational inference to capture this uncertainty in geometry, while BayesRays [10] applies a Bayesian framework to estimate epistemic uncertainty post-training. Our work specifically addresses cleanup of artifacts in static scenes caused by epistemic uncertainty.
@@
Various methods have been proposed to address the visual artifacts common in NeRF reconstructions. Nerfbusters [31] introduced a post-processing technique using diffusion models that learns a local 3D data-driven prior from synthetic data. This data-driven prior, incorporated during NeRF optimization, encourages plausible geometry but requires extensive training and NeRF fine-tuning, adding significant computational overhead. Another approach, BayesRays [10], leverages a post-hoc solution by thresholding an uncertainty field to mask high-uncertainty regions, effectively cleaning artifacts at inference time but at the cost of increased computation during rendering.
These methods, though effective, often come with trade-offs in memory consumption, cleanup time, or inference speed. In contrast, our approach removes artifacts efficiently, preserving the original NeRF’s structure and performance without requiring additional models or extensive post-processing, based only on our Free Space Prior. This ensures both speed and practicality in artifact cleanup.
@@
Data-driven optimization is standard in deep learning, but numerous studies demonstrate that incorporating a well-chosen prior can significantly boost performance. For instance, priors have proven effective in inverse problems like image denoising [5, 12, 33, 27, 6, 7]. In particular, [6, 7] introduce a global image prior enforcing sparsity, showcasing how a Bayesian framework can yield a simple yet powerful denoising algorithm. In the context of 3D reconstruction, priors and regularization play similarly critical roles. Nerfbusters [31] uses a local prior to improve scene fidelity and RegNeRF [24] employs a 2D Total Variation (TV) regularizer to rendered images, both maintaining the standard NeRF architecture as we do. Other methods modify the architecture. For example, Plenoxels [35] employs a 3D Total Variation (TV) regularizer to reduce abrupt density changes between neighboring voxels and 2D TV regularization is also applied in factorized plenoptic fields [4, 9]. Plenoctrees [36] incorporates a sparsity loss within a complex pipeline for faster rendering but at high memory and training costs. To address anti-aliasing, Zip-NeRF [3] and MipNeRF-360 [2] applies distortion loss and interlevel loss, reducing artifacts and improving consistency. While effective, this regularization alone does not eliminate floaters, as our results show, highlighting the need for a stronger constraint on unseen regions.
@@ 3 Background
Understanding NeRF is essential to grasp the significance of our work and its contributions to the field. In this section, we review the fundamental concepts of NeRF and discuss the challenges associated with artifact generation in unobserved regions, which our approach aims to overcome.
@@ 3.1 Neural Radiance Fields
NeRF [22] is an approach for synthesizing novel views of 3D scenes. Each point in 3D space is represented by a view-dependent radiance and a view-independent density, expressed as follows:
@@
Here, $\boldsymbol{\phi}$ denotes the learnable parameters of the neural network. The color along a ray $\mathbf{r}=\mathbf{o}_{\mathbf{r}}+t\cdot\mathbf{d}_{\mathbf{r}}$ is determined by sampling points $t_{i}$} along the ray:
@@
An important observation from Eq.˜2 is that NeRF’s optimization occurs on a per-pixel basis, with the color of each pixel computed along a single ray. Since density decreases exponentially along the ray after it intersects the first surface, there is effectively no optimization beyond that surface, and certainly none where the ray does not pass through. Consequently, unseen regions remain unoptimized, often resulting in unwanted artifacts in these areas.
@@ 4 Method
Our approach, as shown in Fig.˜2, is designed as a post-hoc refinement. It takes a pre-trained NeRF model along with its training cameras and optimize it to remove artifacts in unseen regions while retaining the scene’s intended features.
To date, NeRF models, and NeRF cleanup methods in particular, searched for a density field $\sigma$ that best explains the given data. That is, they look for the maximum likelihood solution to the problem. Some methods extend this by learning a general data distribution over multiple 3D scenes, creating a local prior for artifact removal. In contrast, we apply a simple, global prior on the density $\sigma$ - We take the prior $P(\sigma)$ to be the zero density prior.
This global approach not only simplifies the optimization process but also ensures versatility and robustness across different NeRF architectures and datasets. Unlike learned priors that depend on specific scene distributions, our method operates effectively regardless of scene complexity or data variations, making it widely applicable and independent of dataset-specific assumptions.
Figure 3: Qualitative Cleanup Results: Comparison of cleanup methods on the Plant and Roses scenes from the Nerfbusters dataset, with inference time per frame shown. Each image is a novel view rendered post-cleanup, highlighting the balance between artifact removal and scene coverage. Methods like Free Space Prior and BayesRays (0.9) achieve high coverage with minimal artifacts, while Nerfbusters and BayesRays (0.3) trade coverage for stronger cleanup. Our approach achieves similar results to BayesRays with 40% rendering time.
We base our method on the assumption that “There is nothing in the unseen regions”. While this assumption does not necessarily hold in real-world scenarios, for NeRF reconstructions, it is more practical than allowing unseen regions to contain unregulated noise. To enforce this assumption, we apply a sigmoid-softened density prior term that gently drives density $\sigma$ toward zero.
@@
Ideally, we would like to directly set densities to zero only in the unseen regions. However, determining which regions are strictly unobserved by training rays is challenging due to scene occlusions and structural complexity. Additionally, even if feasible, this direct approach would be computationally expensive. So, instead of explicitly setting densities, we apply the Free Space Prior uniformly across the entire 3D space, affecting both seen and unseen regions. This approach simplifies computation while guiding NeRF optimization to achieve cleaner density distributions.
To implement this, we randomly sample $\mathcal{N}$ points across the 3D space and query the NeRF model $\boldsymbol{\phi}$ for densities at these points, resulting in a set $\Sigma=\{\sigma(\mathbf{x}_{i})\mid\mathbf{x}_{i}\in\mathbb{R}^{3},\,i=1,\dots,N\}$ of density values. We then construct a softened density term for the Free Space Prior, which we integrate into a Free Space Prior Loss term:
$\mathcal{L}_{\text{FSP}}=\sum_{i=1}^{N}\left(\frac{1}{1+e^{-\sigma(\mathbf{x}_{i})}}\right)^{2}$ (4)
This Free Space Prior Loss optimizes the NeRF parameters $\boldsymbol{\phi}$ by enforcing densities to be zero, reflecting our assumption that “There is nothing in the unseen regions”. However, this loss also inadvertently reduces densities in visible regions, where densities should ideally reflect the actual scene structure.
Balancing Cleanup and Scene Integrity
Applying the Free Space Prior Loss indiscriminately across 3D space can lead to an empty scene, as the randomly enforced zero-density constraint propagates, causing the MLP to approximate densities as zero throughout the space. To counter this, we introduce a combination of two competing losses: the Free Space Prior Loss Eq.˜4 to reduce densities in unseen regions and the NeRF photometric loss Eq.˜3 applied along training rays to preserve the scene structure and appearance.
@@
  • Unseen Regions: Only the Free Space Prior Loss affects these regions, pushing densities toward zero and removing floaters. In particular, the densities in occluded regions are also set to zero.
@@
The overall loss function is:
$\mathcal{L}=\mathcal{L}_{\text{rec}}+\lambda\mathcal{L}_{\text{FSP}}$ (5)
where $\lambda$ controls the trade-off between preserving scene structure and removing artifacts. We set $\lambda=0.1$ for all visualizations in the paper. $\lambda$ analysis is in the supplementary A.5.
4.1 NeRF Regularization
NeRF reconstruction is inherently an ill-posed problem — many solutions can fit the observed images. The highly expressive MLP satisfies the photometric loss along training rays while potentially assigning arbitrary densities in unobserved regions. This results in an under-constrained system where only a subset of solutions is physically meaningful. Drawing a parallel to Ridge regression, we consider $L_{2}$ regularized MLP to represent the NeRF. The $L_{2}$ regularizer directly penalizes the parameters, favoring the minimum-norm solution and stabilizing an ill-posed problem.
Yet, there is a difference between the two. $L_{2}$ regularization places a prior on the weight of the MLP, whereas our Free Space Prior places a prior on the densities themselves. Specifically, instead of regularizing the network weights directly, we regularize a specific component of the output — the predicted densities. Since our objective is formulated as a loss optimized via backpropagation, we explicitly sample the 3D space to enforce this regularization in unobserved regions. This modified approach resolves the ill-posedness by favoring solutions with minimal densities in free space, while preserving the correct photometric reconstruction in observed regions.
Figure 4: Density Along a Ray: (Left) Novel view rendered with the Nerfacto Baseline, with a red dot marking the sampled ray and the density along the marked ray. (Right) The same novel view after our cleanup. Before cleanup, the density along the ray is noisy, with floaters causing unintended peaks. After cleanup, only the surface is captured, and densities in unseen regions are enforced to zero. Densities are sigmoid softened and shifted to [0,1].